### v2

@@ Results
The next example is the Zachary Karate Club network [39], a well-known dataset encoding social interactions. The current agreement, for this system describes 4 communities with a maximum modularity of 0.44904 for $\gamma=1$ (Table 1). As opposed to the clique chains studied before, the topological and geometrical structure of this network does not suggest a well-defined hierarchical structure, thus not explicitly favoring the recursive nature of our solution. We also tested a recent quantum-classical hybrid [25], known to correctly resolve this network. Noteworthy, the network is undirected and weighted, and the solution obtained through pure Hierarchical quantum annealing is entirely equivalent.
Algorithm Max $\mathbf{Q}$ Frequency (%) Number of communities Time (s) [mean $\pm$ SEM]
H. annealing 0.444904 100 4 8.46 $\pm$ 0.26
@@
Table 1: Modularity of the Karate club network. Maximum modularity resulting from the different algorithms tested. The maximum and frequencies are taken from a pool of 50 independent runs (see Methods). For the quantum-based solvers, the times include the embedding of the QUBO, its caching, and the communication with the D-Wave servers. In italics, we highlight the methods that directly stem from our algorithm.
Hierarchical annealing as a function of network topology
@@
We studied the performance of the proposed annealing process as we varied the geometry and the topology of the networks. We evaluated the solutions of the Louvain, Leiden, Bayan, and hierarchical annealing algorithms for 3 different complex types of networks and number of nodes. The quantum-classical hybrid algorithm was omitted due to its high consumption of computing minutes. A similar argument holds for Hierarchical Gurobi and Bayan, where external computational resources were needed to be acquired.
@@
The first networks were generated following preferential attachment and triad formation steps [35]. This generates a network with scale-free and/or power-law properties, high clustering, and a good modular organization as the size of the network is increased. The number of edges to add and the probability of performing a triad formation step were kept at 1 and 0.1 respectively for all the sizes. Hierarchical annealing displayed a robust, optimal, and highly congruent behavior when compared to the other algorithms (Fig. 2a).
@@
Then, we generated networks using only preferential attachment (i.e., without triad formation) [34]). However, depending on the number of edges added at each step, the final network may or may not display a modular organization. To test the robustness of hierarchical annealing to discover communities in potentially modular networks, we kept the aforementioned parameter equal to 40% of the network size. This generated poorly modular systems but without random topologies. For these non-trivial networks, our method remained competitive across sizes largely surpassing the Leiden algorithm (Fig. 2b).
@@
We also tested the hierarchical annealing process in random networks [33], where the probability of two nodes being connected is uniform. The topology of Erdős-Rényi networks has been well characterized and is considered a valid null system for multiple graphs and topological metrics in biological systems. Again, our solution produced comparably optimal solutions for all algorithms (Fig. 2c).
@@
Finally, we studied the performance of the hierarchical quantum annealing solution in directed scale-free networks generated using the preferential attachment mechanism [36]. Again, our solution resulted in high-quality community structures for multiple network sizes (Fig. 2d), and even in important increments in certain cases w.r.t. the classical benchmarks (Fig. 2e).
Figure 2: Annealing solutions for different random networks of increasing size. Modularity of a random network as found by the Louvain, Leiden, and Hierarchical annealing algorithms. The resolution, $\gamma$, was set to 1. To measure the overlap between the different solutions, we show the Dice score between all the communities found. The colormaps indicate if the annealing solution has a higher (in blue), equal (in black), or lower (in red) modularity. The opacity is directly proportional to the magnitude of the Dice score between each pair of communities. The random networks have been generated via the mechanisms of (a) power-law clustering, (b) preferential attachment, and (c) random edge creation. d Exact same layout for a directed scale-free network. e Community structure of a directed scale-free network of 60 nodes discovered through the H. annealing and Louvain algorithms. This corresponds to the network where the annealing solution was significantly superior in d.
@@
To further study the robustness of the quantum solutions, we studied the behavior of the hierarchical annealing process as a function of the power-law exponent. For that, we generated multiple networks of different sizes following the preferential attachment mechanism described previously, each one with significantly different power-law structures, and potential hierarchies (see Methods). We compared the hierarchical annealing process against the Louvain and Leiden algorithms finding very similar solutions in terms of the modularity index. Even more, the solutions showed the same inverse relationship on the power-law structure as the classical counterparts (Fig. 3a).
@@
More specifically, the hierarchical annealing process yielded 3 out of 36 equally optimal solutions, 16 out of 36 higher modularity scores, and 13 out of 36 worse solutions (Fig. 3b). We found that the quantum alternative tended to return more optimal solutions for networks with medium-big size and smaller power-law exponents. Therefore, we fitted a bilinear model using an ordinary least squares procedure to explain the relative increase (%$(\%RI)$ of the quantum solution as a function of both parameters,
%$\%RI\sim 1+\alpha+N,$
@@
where $\alpha$ denotes the power-law exponent, and $N$ is the number of nodes. This model carried explanatory power ($R^{2}=0.3680$, $p=0.0080$, $F$-statistic against a constant model, AIC=118.2572) and both coefficients significantly contributed to the prediction ($\alpha$: $p=0.0.0116$, $N$: $p=0.0462$). Additionally, we included an interaction term of the form $\alpha N$. This alternative model showed a slightly higher coefficient of determination ($R^{2}=0.3869$, $p=0.0144$, $F$-statistic against a constant model, AIC=119.5288) but the individual coefficients were non-significant ($\alpha$: $p=0.0863$, $N$: $p=0.7353$, $\alpha N$: $p=0.4416$), suggesting that independent rather than interaction effects better explain the performance of the algorithm.
@@
To visualize and confirm these trends, we fitted two independent linear models for each parameter. The coefficient of the power-law exponent maintained the significance, while the coefficient of the number of nodes didn’t despite the increasing trend in the relative increase (Fig. 3c). For scale-free networks (i.e., $\alpha\in[2,3]$), all the hierarchical annealing solutions were either superior or equivalent to ones found by the classical counterparts (Fig. 3c, green shaded area). Importantly, we only considered solutions with a relative increase and power-law exponent that fell within the 95% range (i.e., $|z_{score}|<1.96$). Using this criteria, 4 solutions out of 36 were discarded. Crucially, 3 out of these 4 outliers, corresponded to extremely high power-law exponents.
@@
Figure 3: Hierarchical annealing as a function of inherent hierarchies in the network. a Modularity of the solution for the Leiden, Louvain, and Hierarchical annealing algorithms as a function of the number of nodes and scale-freeness of the networks. The networks were generated using a preferential attachment mechanism [34]. b Explicit comparison between the Hierarchical annealing and Leiden (top) and Louvain (bottom) solutions on the plane. Black markers (located on the dotted diagonal line) correspond to identical solutions, blue markers correspond to better solutions from the annealer, and red markers indicate worse performances than the classical counterparts. The shape of the markers follows the top-right legend in panel a. c Relative increase of the quantum solutions in a, b as a function of $\alpha$, and the number of nodes. The green shaded area (top-left) indicates scale-free networks. The solutions in black markers (bottom) were discarded for the regression analyses.
Hierarchical annealing for different resolution parameters
@@
The binary hierarchical formalism we developed in the Methods, explicitly incorporated the resolution parameter $\gamma$ in the function to optimize, as opposed to earlier accounts, classical [3, 28] and quantum [23, 24]. Briefly, the process was formally independent of $\gamma$, which only appeared in the calculation of the modularity matrix (see Methods). We generated 3 different networks with nodes $N=10,50,80$ using preferential attachment and triad formation steps to test whether the framework returned an optimal community structure. We then ran our Hierarchical annealing procedure for different values of the resolution parameter.
The overlap between the different solutions was consistently close to 1, and never below 0.8, as measured with the average Dice score between all the communities (Fig. 4a). This overlap, as expected, co-varied with the number of communities found. Furthermore, the performance was stable across the range of $\gamma$ studied, showing some slight increases and decreases for individual cases (Fig. 4b-c).
We kept track of the total computing time for the hierarchical process to finish (see Methods). This time increased rather monotonously with the resolution (Fig. 4c). However, larger values of $\gamma$ are related to smaller communities [31], thus the number of calls to the QPU also increased (Fig. 4).
@@
Figure 4: Hierarchical annealing in different clustered power-law networks for different resolution parameters. Each column depicts the results described below for a network of $N=10,50,80$ nodes respectively. a Measure of the average overlap between the communities found by the Louvain and Hierarchical annealing algorithms (left axis, blue) and the number of communities (right axis, salmon and black) as a function of the resolution parameter $\gamma$. b Relative increase of the Hierarchical annealing measured w.r.t. the Louvain (top) and Leiden (bottom) solutions in a. Black markers correspond to identical solutions, blue markers correspond to better solutions from the annealer, and red markers indicate worse performances than the Louvain alternative. c Maximum modularity per resolution value (left axis, same legend as in a) and the total computing time per solution (right axis, green).
Application of QA to structural brain connectivity
@@
One of the main advantages of our proposal is that we could inspect the output process at every step of the process. We plotted the modularity and the community structure in the form of a dendrogram. The final division was dependent on each earlier subdivision, thus being interpretable as a hidden community structure. Such hidden structures would be informative of certain network vulnerabilities or strengths. We applied this visualization technique to all the networks studied up until now, while comparing the modularity of the final community structure with the same benchmark methods achieving competitive and identical results (Fig. S.3; see also Supplementary Figures).
@@
Figure 5: Dendrogram of a clustered power-law network [35] of N=100 nodes. Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 15 and resolution = 1. Both the Louvain and Leiden algorithms returned a maximum modularity of $Q=0.7779$. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments. The bottom-most row represents the final output of the H. annealing algorithm.
@@
As a last proof of concept, we analyzed real brain connectivity networks as derived from diffusion-weighted magnetic resonance imaging methods. These networks contained a total of 166 nodes and were densely connected while exhibiting somewhat scale-free structures [40]. We also inspected the anatomical positioning and organization of the found community structure as well as the hierarchical levels of the process (Fig. 6). Despite the large number of variables (i.e., qubits), the H. annealing solution was strikingly close to the solutions found by the Louvain and Leiden algorithms. Furthermore, the communities found by the quantum algorithm only differed slightly in the frontoparietal border, areas that are highly interconnected with U-fibers. The rest of the identified communities were largely in agreement between algorithms (see Supplementary Figures), hence setting a very promising set of baseline results for more advanced and sophisticated QA methods and algorithms.
@@
Figure 6: Dendrogram of a real brain structural connectivity network [40]. a Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 10 and resolution = 1. The Louvain and Leiden algorithms returned a maximum modularity of $Q=0.611$ and $Q=0.610$ respectively. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments (i.e., Eq. 9). The bottom-most row represents the final output of the H. annealing algorithm. b Overlay of the structural communities on the Montreal Neurological Institute (MNI) 152 template. Each color corresponds to a different community, whose elements can be found in a.
Discussion

### v3

@@ 3.1 Hierarchical annealing in simple networks
The next example was the Zachary Karate Club network, a well-known dataset encoding social interactions. The current agreement for this system describes 4 communities with a maximum modularity of 0.44904 for $\gamma=1$ (Table 2). As opposed to the clique chains studied before, the topological and geometrical structure of this network does not suggest a well-defined hierarchical structure, thus not explicitly favoring the recursive nature of our solution. We also tested a recent quantum-classical hybrid [Wierzbinski2023], known to correctly resolve this network. The solution obtained through pure Hierarchical quantum annealing was entirely equivalent.
Algorithm Max $\mathbf{Q}$ Frequency (%) Communities Time (s) [mean $\pm$ SEM]
H. annealing 0.444904 100 4 n.a.
@@
Spectral clustering 0.444904 100 4 0.00436 $\pm 2\cdot 10^{-5}$
Simulated annealing 0.444904 2 4 3.1772 $\pm$ 0.01731
Infomap 0.434521 100 3 0.000664 $\pm 7\cdot 10^{-6}$
SBM Inference 0.434521 0.01 3 0.0028 $\pm$ 0.0004
@@
Table 2: Modularity of the Karate club network. Maximum modularity resulting from the different algorithms tested. The maximum and frequencies are taken from a pool of 50 independent runs for all algorithms except for SBM inference, which required 1000 runs. In italics, methods that stem from our algorithm. ’n.a.’ stands for not applicable since the computing times are analyzed separately.
3.2 Hierarchical annealing as a function of network topology
@@
We studied the performance of the proposed annealing process as we varied the geometry and the topology of the networks. We evaluated the solutions of the Louvain, Leiden, simulated annealing, spectral clustering, and Hierarchical annealing algorithms for 3 different complex types of networks of increasing sizes. The DQM algorithm was ran in a subset of networks due to its high consumption of computing time. Specifically, for these experiments, DQM required approximately 300× more billed computing time than the Hierarchical annealing procedure, substantially impacting access to computational resources and the planning of computational grants (Table S1). A similar argument holds for Hierarchical Gurobi and Bayan, where external computational resources needed to be acquired. Infomap and SBM inference were not considered here because they do not maximize modularity, but instead discover communities from different principles [Rosvall2008, Peixoto2014].
@@
The first networks were generated following preferential attachment and triad formation steps [Holme2002]. This generated a network with scale-free and/or power-law properties, high clustering, and a good modular organization as the size of the network is increased. The number of edges to add and the probability of performing a triad formation step were kept at 1 and 0.1, respectively, for all the sizes. Hierarchical annealing displayed a robust, optimal, and highly congruent behavior when compared to the other algorithms (Fig. 2a-b).
@@
Then, we generated networks using only preferential attachment (i.e., without triad formation) [Barabasi1999]. However, depending on the number of edges added at each step, the final network may or may not display a modular organization. To test the robustness of Hierarchical annealing to discover communities in potentially modular networks, we kept the aforementioned parameter equal to 40% of the network size. This generated poorly modular systems, but without random topologies. For these non-trivial networks, our method remained competitive across sizes, largely surpassing the Leiden algorithm (Fig. 2c-d). DQM and simulated annealing were robust but substantially more computationally intensive than the other methods. Moreover, simulated annealing required a pre-specified number of communities inferred by Hierarchical annealing, since when initialized at the resolution limit, its performance failed to pose any real challenge to the alternative solutions.
@@
We also tested the Hierarchical annealing process in random networks [erdds1959random], where the probability of two nodes being connected is uniform. The topology of Erdős-Rényi networks has been well characterized and is considered a valid null system for multiple graphs and topological metrics in biological systems. Again, our solution produced comparably optimal solutions for all algorithms (Fig. 2e-f). Similarly to the Barabasi-Albert systems, the DQM and simulated annealing solutions were competitive at the expense of computational tractability and the need for a priori sets of communities to detect.
@@
Finally, we studied the performance of the hierarchical quantum annealing solution in directed scale-free networks generated using the preferential attachment mechanism [bollobas2003]. Again, our solution resulted in high-quality community structures for multiple network sizes (Fig. 2g-j). Both DQM and simulated annealing could resolve these systems, but with a slight underperformance.
Figure 2: Hierarchical annealing solutions in random networks. a Maximum modularity ($\max Q$) of power-law–clustered networks as a function of network size. b Consensus matrices for the $N=100$ network in (a), comparing communities identified by the hierarchical annealing procedure with those obtained using alternative algorithms. Opacity is proportional to the Dice similarity score between pairs of communities. Black, blue, and red outlines indicate cases in which hierarchical annealing achieved equal, higher, or lower modularity, respectively, relative to the corresponding alternative solution. White entries denote zero Dice similarity. c Maximum modularity ($\max Q$) of Barabasi-Albert networks. d Consensus matrices of the Barabasi-Albert network ($N=100$) studied in (c). e Maximum modularity ($\max Q$) of Erdős-Renyi networks. f Consensus matrices of the Erdős-Renyi network ($N=100$) studied in (e). g Maximum modularity ($\max Q$) of directed scale-free. h Consensus matrices of the directed scale-free network ($N=70$) studied in (g). i-j Communities found the Hierarchical annealing (i) and Louvain (j) algorithms in the directed scale-free network in (h). See Fig. S5-S8 for the full set of consensus matrices.
@@
To further study the robustness of the quantum solutions, we studied the behavior of the Hierarchical annealing process as a function of the power-law exponent. For that, we generated multiple networks of different sizes following the preferential attachment mechanism described previously, each one with significantly different power-law structures and potential hierarchies. We compared the Hierarchical annealing process against the Louvain and Leiden algorithms, finding very similar solutions in terms of the modularity index. Even more, the solutions showed the same inverse relationship on the power-law structure as the classical counterparts (Fig. 3a).
@@
More specifically, the Hierarchical annealing process yielded 3 out of 36 equally optimal solutions, 16 out of 36 higher modularity scores, and 13 out of 36 worse solutions (Fig. 3b). Therefore, we fitted a bilinear model using an ordinary least squares procedure to explain the relative increase $(\%RI)$ of the quantum solution as a function of both parameters,
$\%RI\sim 1+\alpha+N,$
@@
where $\alpha$ denotes the power-law exponent, and $N$ is the number of nodes. This model carried explanatory power ($R^{2}=0.3680$, $p=0.0080$, $F$-statistic against a constant model, AIC=118.2572) and both coefficients significantly contributed to the prediction ($\alpha$: $p=0.0.0116$, NN: $p=0.0462$). Additionally, we included an interaction term of the form $\alpha N$. This alternative model showed a slightly higher coefficient of determination ($R^{2}=0.3869$, $p=0.0144$, $F$-statistic against a constant model, AIC=119.5288) but the individual coefficients were non-significant ($\alpha$: $p=0.0863$, NN: $p=0.7353$, $\alpha N$: $p=0.4416$), suggesting that independent rather than interaction effects better explain the performance of the algorithm.
@@
To visualize and confirm these trends, we fitted two independent linear models for each parameter. The coefficient of the power-law exponent maintained the significance, while the coefficient of the number of nodes didn’t, despite the increasing trend in the relative increase (Fig. 3c). For scale-free networks (i.e., $\alpha\in[2,3]$), all the Hierarchical annealing solutions were either superior or equivalent to ones found by the classical counterparts (Fig. 3c, green shaded area). Importantly, we only considered solutions with a relative increase and power-law exponent that fell within the 95% range (i.e., $|z_{score}|<1.96$). Using this criterion, 4 solutions out of 36 were discarded. Crucially, 3 out of these 4 outliers corresponded to extremely high power-law exponents.
@@
Figure 3: Hierarchical annealing as a function of inherent hierarchies in random networks. a Modularity obtained with the Leiden, Louvain, and Hierarchical Annealing algorithms across networks varying in size (panels: from left to right) and scale-freeness (marker shape: ”Fraction” $f$), generated via preferential attachment [Barabasi1999]. b For networks of increasing size (panels: left to right), comparison of Hierarchical Annealing with Leiden (panels: top) and Louvain (panels: bottom). Black markers located on the diagonal indicate identical modularities; blue/red markers located on the upper/lower triangular parts denote superior/inferior annealing solutions. Marker shapes correspond to the legend in (a). c Relative improvement of annealing solutions in (a-b) as a function of $\alpha$ and network size. The green shaded region (top left) corresponds to scale-free networks. Bottom panels are identical but explicitly show cases excluded from the regression analyses (black markers).
An alternative benchmark to evaluate community detection methods is the stochastic block model (SBM), in which communities are planted a priori and must then be recovered. Across a range of difficulty levels, the Hierarchical annealing procedure performed consistently well across several evaluation metrics (Fig. 4; see also Fig. S9). Importantly, these experiments were not intended to validate the consistency of modularity maximization as a community detection method itself [Bickel2009, Zhao2012, Ghasemian2020], but to establish Hierarchical annealing as a valid approach to recover known structure by attaining $\max Q$ [Newman2016].
Figure 4: Hierarchical annealing in stochastic block models (SBM). Performance of benchmark methods relative to the hierarchical annealing procedure on an SBM with three communities ($p_{in}=0.9$, $p_{out}=0.05$). Recovery accuracy is assessed using four metrics against the ground truth, with the corresponding maximum modularity ($\max Q$) shown in the rightmost panel as a function of the mixing parameter $\lambda$.
3.3 Hierarchical annealing for different resolution parameters
@@
The binary hierarchical formalism we developed explicitly incorporated the resolution parameter $\gamma$ in the function to optimize, as opposed to earlier accounts, both classical [Newman2006, Leicht2008] and quantum [Ushijima-Mwesigwa2017, Negre2020]. Briefly, the process was formally independent of $\gamma$, which only appeared in the calculation of the generalized modularity matrix (Methods). However, since $\gamma$ determines both the QUBO coefficients, recursion depth, and physical constraints, we examined its influence.
We generated 3 different networks with nodes $N=10,50,80$ using preferential attachment and triad formation steps to test whether the framework returned an optimal community structure. We then ran our Hierarchical annealing procedure for different values of the resolution parameter and compared it to the Leiden and Louvain alternatives since they were the fastest and most competitive while incorporating $\gamma$ in their formalism. The overlap between the different solutions was consistently close to 1, and never below 0.8, as measured with the average Dice score between all the communities (Fig. 5a). This overlap, as expected, co-varied with the number of communities found. Furthermore, the performance was stable across the range of $\gamma$ studied, showing some slight increases and decreases for individual cases (Fig. 5b-c). A similar situation held for both Erdős-Renyi and Barabasi-Albert networks (Figs. S10-S12), with some slight but still comparable underperformance for $\gamma>2$. In contrast, for directed scale-free networks, the empirical performance of the hierarchical annealing algorithm degraded both for low and high resolution values as the size of the networks grew (Fig. S13).
@@
Figure 5: Hierarchical annealing in different clustered power-law networks for different resolution parameters. Each column depicts the results described below for a network of $N=10,50,80$ nodes, respectively. a Measure of the average overlap between the communities found by the Louvain and Hierarchical annealing algorithms (left axis, blue) and the number of communities (right axis, salmon and black) as a function of the resolution parameter $\gamma$. b Relative increase of the Hierarchical annealing measured w.r.t. the Louvain (top) and Leiden (bottom) solutions in (a). Black markers correspond to identical solutions, blue markers correspond to better solutions from the annealer, and red markers indicate worse performances than the Louvain alternative. c Maximum modularity per resolution value.
3.4 Hierarchical annealing: running times
We measured the computing times of the hierarchical annealing (Algorithm 1) for the same types of networks (Methods). The number of nodes was gradually increased from 10 to 140 to assess how running times scale with problem size. For Power-law Clustered, Albert-Barabási, and Erdős-Rényi networks, experimental times were recorded for a reduced set of resolution parameters and varying sizes. For the Karate Club network, the modularity index was maximized across multiple resolution parameters. As expected, QPU access times increased with the number of nodes for all networks (Fig. 6). For $\gamma=1$, we compared linear and logarithmic models to characterize the scaling, finding that the latter yielded slightly better agreements with empirical times (Table 3). Higher resolution parameters increased QPU access times (see Tables S2-S7).
In Power-law clustered networks, which exhibited inherent hierarchical structure, the effect of $\gamma$ was modest and scaled similarly to the $\gamma=1$ case (Fig. 6a). In contrast, for Barabási–Albert and Erdős–Rényi networks, where no clear hierarchies existed, QPU access times behaved nontrivially, suggesting a stronger influence of $\gamma$ on the problem Hamiltonian. For the Karate Club network, total QPU access times ranged from 0.1062 (s) for $\gamma=0.5$ to 0.3771 (s) for $\gamma=2$, with an approximately linear dependence (Table S8).
Network type Logarithmic $t\sim\log n$ Linear $t\sim n$
Powerlaw clustered $R^{2}=0.96$ $R^{2}=0.94$
Barabasi-Albert $R^{2}=0.91$ $R^{2}=0.87$
Erdős-Renyi $R^{2}=0.82$ $R^{2}=0.58$
Table 3: Determination coefficient $\mathbf{R^{2}}$ for QPU access time fits as a function of network size. Logarithmic and linear fits were used to model $\mathcal{K}_{n}$ (see also Fig. 6).
Cache-read embedding times showed no clear functional dependence on the number of nodes, with variations likely due to internal processes. For Power-law networks, the resolution parameter had no effect (Fig. 6b), while in Barabási–Albert and Erdős–Rényi networks, higher resolutions generally increased reading times (Fig. 6e,h). In contrast, for the Karate Club network, cache-read embedding times ranged from 0.2072 (s) for $\gamma=0.5$ to 3.7889 (s) for $\gamma=2$, indicating a dependence on the specific QUBO values and resulting embeddings.
Interestingly, communication and queueing times represented a substantial overhead relative to actual QPU processing, highlighting inefficiencies that could be mitigated by local annealers or alternative queueing schemes. For Power-law clustered networks, these times consistently increased (Fig. 6c), while they remained stable for Barabási–Albert and Erdős–Rényi networks (Fig. 6f,i). The effect of the resolution parameter was inconsistent, likely due to the larger magnitude and unclear dependence of cached-embedding read times compared to QPU access times. For the Karate Club network, total sampling and communication times ranged from 7.1366 (s) for $\gamma=0.5$ to 30.8441 (s) for $\gamma=2$.
Figure 6: Experimental measurements of QPU access, embeddings, and Sample times. a The total QPU access times as an empirical measure of $\mathcal{K}_{n}$ in Eq. 18, corresponding to the sum of the repeated steps of binary QA optimization in power-law clustered networks of increasing size and different resolution parameters, $\gamma$. The inset depicts linear fits between measured times and the natural logarithm of network size ($t\propto\ln N$). b The total embedding and reading times associated with mapping the power-law networks in (a) to the QPU topology in the Advantage system. c The total time required to execute the DWAVE Sample function for the power-law networks in (a). d-f Identical to (a-c) but in Barabasi-Albert networks. g-i Identical to (a-c) but in Erdős -Renyi networks. See Supplementary Material for the exact numbers corresponding to the results shown here.
3.5 Application of QA to structural brain connectivity
@@
One of the main advantages of our proposal is that we could inspect the output process at every step of the process. We plotted the modularity and the community structure in the form of a dendrogram. The final division was dependent on each earlier subdivision, thus being interpretable as a hidden community structure. Such hidden structures would be informative of certain network vulnerabilities or strengths. We applied this visualization technique to all the networks studied up until now, comparing the modularity of the final community structure with that of the same benchmark methods, achieving competitive and identical results (Fig. 7; see also Fig. S14-S18).
@@
Figure 7: Dendrogram of a clustered power-law network [Holme2002] of N=100 nodes. Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 15 and resolution = 1. Both the Louvain and Leiden algorithms returned a maximum modularity of $Q=0.7779$. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments. The bottom-most row represents the final output of the H. annealing algorithm.
@@
As a last proof of concept, we analyzed real brain connectivity networks as derived from diffusion-weighted magnetic resonance imaging methods. These networks contained a total of 166 nodes and were densely connected while exhibiting somewhat scale-free structures [falco2024functional]. We also inspected the anatomical positioning and organization of the found community structure as well as the hierarchical levels of the process (Fig. 8). Despite the large number of variables (i.e., qubits), the H. annealing solution was strikingly close to the solutions found by the Louvain and Leiden algorithms. Furthermore, the communities found by the quantum algorithm only differed slightly in the frontoparietal border, areas that are highly interconnected with U-fibers. The rest of the identified communities were largely in agreement between algorithms (Fig. S19-S20), hence setting a very promising set of baseline results for more advanced and sophisticated QA methods and algorithms.
@@
Figure 8: Dendrogram of a real brain structural connectivity network [falco2024functional]. a Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, $N_{runs}=10$ and $\gamma=1$. The Louvain and Leiden algorithms returned a maximum modularity of $Q=0.611$ and $Q=0.610$, respectively. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments. b Overlay of the structural communities on the Montreal Neurological Institute (MNI) 152 template. Colors correspond to the communities in (a).
4 Discussion
Here we present a quantum approach to discovering communities in complex networks. We unify the optimization process for binary, weighted, and directed networks and explicitly consider the resolution parameter to consider multiple topological scales present in the networks. Our experiments show that quantum annealing (QA) can be successfully used to maximize the modularity of a network and pinpoint meaningful groups of highly connected nodes. Its performance is robust across multiple experimental conditions and, at least comparable to alternative methods based on classical heuristics. However, the recursion in Algorithm 1 does not guarantee global optimality, nor do quantum annealing, simulated annealing, or other heuristics considered here for this NP-hard problem. From a theoretical standpoint, quantum and simulated annealing are closely related: both are probabilistic methods that sample Boltzmann-like distributions, with quantum fluctuations in QA playing a role analogous to thermal fluctuations in simulated annealing. While prior works study problem-dependent advantages of QA without universal superiority [Somma2007, Nishimori2015, Wauters2017], results on current hardware remain mixed due to noise and finite annealing effects [Okada2019]. In this context, our results show that both simulated and hierarchical QA achieve competitive performance, with the latter offering practical advantages by avoiding an a priori choice of the number of communities. Simulations of quantum hardware and processes could help assess whether any advantage persists at larger scales [King2021], but they are computationally expensive and currently limited to systems with only a few physical qubits [Chen2022, morrell2024quantumannealing].